8-th International Workshop on Complex structures and vector fields

Sofia, 21-26 August 2006

PROGRAM

August 18, 2006

AUGUST 21 Differential Geometry

8:00 – 9:15  REGISTRATION
9:15 – 9:45  OPENING

(Dodunekov, J. Stamenov, S. Dimiev, V. Gerdjikov)

9:45 – 10:30 Yasuo Matsushita Almost Kähler-Einstein structures on 8dimensional Walker manifolds

10:30 – 10:50 Coffee break

10:50 - 11:35 Peter Popivanov On the hypoellipticity of complex valued vector

fields

11:35 - 12:15 Toshiaki Adachi Killing helices on a complex projective space and canonical magnetic fields on geodesic spheres

Lunch time

15:30 - 16:10 Vestislav Apostolov Generalized Kähler manifolds, commuting

complex structures and split tangent bundles

16:10 - 16:50 Rossen Dandoloff Anholonomy of a moving space curve and the Schrdinger equation

16:50 - 17:10 Coffee break

17:10 - 17:50 Hideya Hashimoto On almost Hermitian structures of 6-

dimensional submanifolds in the octonions

17:50 - 18:30 Norio Ejiri Complex submanifolds and Lagrangian submanifolds associated with minimal surfaces in tori

AUGUST 22 Mathematical Physics and PDE

9:00 - 9:45 Julian Lawrinowicz Clusters and fractals in mathematical

physics

9:45 - 10:30 Vladimir Georgiev Nonexistence of nonzero resonances for Schrdinger operators with singular perturbation

10:30 10:50 Coffee break

10:50 - 11:30 Vladimir Gerdjikov Hamiltonian aspects of soliton equations

with deep reductions

11:30 - 12:00 Tihomir Valchev Bäcklund transformations and Riemann-Hilbert problem for N-wave equations with additional symmetries

Lunch time

15:30 - 16:10 Georgi Grahovski The Caudrey-Beals-Coifman systems and the

gauge group action

16:10 - 16:50 Nikolay Kostov The Manakov model as two moving interacting curves

16:50 - 17:10 Coffee break

17.10 - 17.50 Stoil Donev Integrability, curvature and description of photon-like

objects

17:50 - 18:30 Tadashi Sugiyama Characterization of totally umgilic immersions by curves of order 2

AUGUST 23 Differential Geometry

9:00 - 9:45  Kouei Sekigawa Some critical almost Kähler structures

9:45 - 10:25 Sadahiro Maeda Characterization of parallel isometric immersions of space forms into space forms in the class of isotropic immersions

10:25 - 10:45 Coffee break

10:45 - 11:15 Milen Hristov On the real hypersurfaces of Hermitian manifolds

11:15 - 11:45 Galya Nakova On some non-integrable almost contact manifolds

with Norden metric of dimension 5

11:45 - 12:25 Bozhidar Iliev Normal frames and linear transports along paths in line bundles. Applications to classical electrodynamics

Lunch time 15:30 - 16:00 Velichka Milousheva On the geometric structure of hypersur-

faces of conullity two in Euclidean space

16:00 - 16:40 Mancho Manev On quasi-Kähler manifolds with Norden metric

16:40 - 17:00 Coffee break

17:00 - 17:30 Dobrinka Gribacheva Lie groups as four-dimensional Rieman-

nian product manifolds

17:30 - 18:00 Marta Teofilova Lie groups as several complex manifolds with Norden metric

Conference dinner

AUGUST 24 EXCURSION

AUGUST 25 Complex Analysis

9:00 - 9:40 Vladimir Balan KCC and linear stability for the Parkinson

tremor model

9:40 - 10:20 Stancho Dimiev Bi-complex analytic pseudo-Euclidean geometry and applications

10:20 - 10:40 Coffee break

10:40 - 11:10 Nikolai Nikolov On the definition of Kobayashi-Buseman pseu-

dometric

11:10 - 11:40 Petar Stoev Differential forms of bi-complex variables
11:40 - 12:10 Petya Furlinska Polynomials of bi-complex variables

Lunch time

15:30 - 16:00 Branimir Kiradjiev Quadratic forms on the bi-complex algebras

16:00 - 16:30 Lilya Apostolova Real analyticity of the hyper-Kähler almost Kähler manifolds

16:30 - 16:50 Coffee break

16:50 - 17:20 Assen Kyuldjiev Symmetries of the Manev problem and its real

Hamiltonian form

17:20 - 18:00 Rossen Ivanov Camassa-Holm equation as a geodesic flow of the right invariant metric

AUGUST 26 Complex Analysis and Applied Mathematics

9:00 - 9:40 Hiroshi Matsuzoe Information geometry and affine differential

geometry

9:40 - 10:20 Vesna Manova Distributional boundary values of the functions in the Djarbjashian classes

10:20 - 10:40 Coffee break

10:40 - 11:20 Mihail Konstantinov Effects of finite arithmetic – myths and

realities

11:20 - 11:50 Angel Ivanov, George Venkov Global existence of the solution

to the Hartree equation

11:50 - 12:20 Ilona Zasada Phase transitions in binary alloy thin films

12:20 - 12:50 Marin Marinov Homogenization Of Nonlinear Parabolic Operators Of High Order

Lunch time

15:30 - 16:00 Vanya Markova Application of Unscented and Extended Kalman

Filtering for Estimating Quaternion Motion

16:00 - 16:30 Peace Nwaobilor Non-oscillating trajectories of vector fields and

Zariskis local uniformization

16:30 - 17:00 Clement Ijeh Centroidal Voronoi tessellation based algorithms for vector fields visualization and segmentation

17:00 - 17:20 Coffee break

17:20 - 17:50 Oladipupo Ipadeola Anisotropic depth migration on complex-

structure land data

17:50 - 18:20 Isaak Ibenme Dislocations in complex materials
18:20 - 18:50 CLOSING